The sum of the length and width of a rectangle is 29. The length is 15 less than 3 times the width. Find the dimensions of the rectangle.
Answer by ReadingBoosters(3246)

l + w = 29
l = 3w - 15
Substitute the second equation into the first equation
(3w-15) + w = 29
Combine like terms
4w = 29 + 15
4w = 44
w = 11
l = 3(11) - 15 = 33-15 = 18